Frank can type a report in 3 hours and James takes 7 hours. How long will it take the two of them typing together?
Answer by stanbon(75887)

Frank can type a report in 3 hours and James takes 7 hours. How long will it take the two of them typing together?
----------------------
Frank rate = 1/3 job/hr
James rate = 1/7 job/hr
Together rate = 1/x job/hr
------------------
Equation:
rate + rate = together rate
1/3 + 1/7 = 1/x
(7+3)/21 = 1/x
x = 21/10
x = 2.1 hrs = 2 hr 6 minutes (time to do the job together0
=============
Cheers,
Stan H.